An aircraft with a mass of 936 kg accelerates from 0 to 94 m/s in 9.1 seconds, at which time the aircraft launches. What is the kinetic energy of the aircraft when the aircraft launches?
Answer #1

1) v=u+at

94=0+a*9.1
a=10.33

F=m*a=936*10.33=9668.57

v^2-u^2= 2as

s=427.68m

workdone = kinetic energy = Fs= 9668.7*427.68=4135171.016